OpenStudy (anonymous):

use the formula to evaluate the series 1-2+4-8+ ...a^10

OpenStudy (anonymous):

please help

OpenStudy (jdoe0001):

\(\begin{array}{rrllll} 1&\\ -2&\implies 1\cdot -2\\ 4&\implies -2\cdot -2\\ -8&\implies 4\cdot -2 \end{array}\) what do you think?

OpenStudy (anonymous):

-2x-4x-8

OpenStudy (jdoe0001):

well.. notice that the next term is just the previous one multiplied by -2

OpenStudy (anonymous):

so -8x2 then 16x2

OpenStudy (jdoe0001):

notice that the next term is just the previous one multiplied by \(\bf \color{red}{-2}\)

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

I don't get this. Thanks for helping

OpenStudy (jdoe0001):

the next one will be -8 * -2 = 16 then the next one will be 16 * -2 = -32

OpenStudy (jdoe0001):

\(\large a^{10}\) just means the 10th term in the sequence